Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

radio station

Expert replies
by dreamv » Thu Feb 23, 2012 11:25 pm
On his drive to work, Leo listens to one of three radio stations, A,B,or C. He first turns to A. If A is playing a song he likes, he listens to it;if not, he turns to B. If B is playing a song he likes, he listens to it; if not, he turns to C. If C is paying a song he likes, he listens to it;if not, he turns off the radio. For each station, the probability is 0.30 that at any given moment the station is playing a song Leo likes. On his drive to work, what is the probability that Leo will hear a song he likes?

A. 0.027
B. 0.090
C. 0.417
D. 0.657
E. 0.900
Join the discussion
Source: — Problem Solving |

by kul512 » Fri Feb 24, 2012 2:01 am
Following is the probability-

(He listens songs on A) OR (He listens songs on B) OR (he listens songs on C).

We need to add the probability of all three cases-
1) Probability that he listens song on Channel A= 0.3

2) Probability that he listens song on Channel B=he doesn't listen song on channel AX listen song on channel B
=(1-0.3)X0.3=0.7X0.3=0.21

3) Probability that he listens song on channel C= He doesn't listen song on channel AX he doesn't listen song on channel B X Listen song on channel C
=(1-0.3)X(1-0.3)X0.3=0.49X0.3=0.147

SO total=
[spoiler]0.3+0.21+0.147=0.657

Choose (D)[/spoiler]
Join the discussion

by Brent@GMATPrepNow » Fri Feb 24, 2012 6:53 am
dreamv wrote:On his drive to work, Leo listens to one of three radio stations, A,B,or C. He first turns to A. If A is playing a song he likes, he listens to it;if not, he turns to B. If B is playing a song he likes, he listens to it; if not, he turns to C. If C is paying a song he likes, he listens to it;if not, he turns off the radio. For each station, the probability is 0.30 that at any given moment the station is playing a song Leo likes. On his drive to work, what is the probability that Leo will hear a song he likes?

A. 0.027
B. 0.090
C. 0.417
D. 0.657
E. 0.900
Here's another approach.

P(likes a song) = 1 - P(likes no song)

P(likes no song) = P(doesn't like song on channel A AND doesn't like song on channel B AND doesn't like song on channel C)
= P(doesn't like song on channel A) X P(doesn't like song on channel B) X P(doesn't like song on channel C)
= (0.7)(0.7)(0.7)

Aside we can approximate this by recognizing that (0.7)(0.7) is about 0.5
So, (0.7)(0.7)(0.7) ~ (0.5)(0.7) ~ 0.35

So, P(likes no song) = 0.35

This means that P(likes a song) = 1 - 0.35
= 0.65

The closest answer is D

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by kul512 » Fri Feb 24, 2012 5:40 pm
Brent@GMATPrepNow wrote:
dreamv wrote:On his drive to work, Leo listens to one of three radio stations, A,B,or C. He first turns to A. If A is playing a song he likes, he listens to it;if not, he turns to B. If B is playing a song he likes, he listens to it; if not, he turns to C. If C is paying a song he likes, he listens to it;if not, he turns off the radio. For each station, the probability is 0.30 that at any given moment the station is playing a song Leo likes. On his drive to work, what is the probability that Leo will hear a song he likes?

A. 0.027
B. 0.090
C. 0.417
D. 0.657
E. 0.900
Here's another approach.

P(likes a song) = 1 - P(likes no song)

P(likes no song) = P(doesn't like song on channel A AND doesn't like song on channel B AND doesn't like song on channel C)
= P(doesn't like song on channel A) X P(doesn't like song on channel B) X P(doesn't like song on channel C)
= (0.7)(0.7)(0.7)

Aside we can approximate this by recognizing that (0.7)(0.7) is about 0.5
So, (0.7)(0.7)(0.7) ~ (0.5)(0.7) ~ 0.35

So, P(likes no song) = 0.35

This means that P(likes a song) = 1 - 0.35
= 0.65

The closest answer is D

Cheers,
Brent

Definitely this approach is better.
Sometimes there is very fine line between right and wrong: perspective.
Join the discussion